# Multiport Analytical Pixel Electromagnetic Simulator (MAPES) for AI-assisted RFIC and Microwave Circuit Design

### v1

@@ I Introduction
Secondly, for specific specifications (targeted scattering parameters), the developed efficient EM simulators can be utilized for inverse synthesis of EM structures using heuristic algorithms [5], such as genetic algorithms (GA), simulated annealing, or generative AI tools such as auto-encoders, tandem neural networks [15], or diffusion models [12].
@@
Another technique conceptually related to MAPES, the Method of Moments (MoM)-based "efficient calculator" in [27], precomputes the impedance matrix of a "mother" structure and accelerates exploration by directly manipulating this matrix (removing rows/columns) to evaluate substructures, thereby avoiding repeated matrix refills/factorizations. However, that framework was devised for planar plates and thin-wire geometries excited by voltage/current sources, making it difficult to incorporate vertical ports, vias, and lumped or active components-elements that are intrinsic to MW/RFIC pixels. In contrast, MAPES natively captures horizontal and vertical couplings and supports arbitrary loads, including realistic nonzero via impedances and comprehensive I/O modeling. Furthermore, the MAPES prior-data impedance matrix can be obtained directly from commercial tools (e.g., CST, HFSS), while the mesh-level MoM impedance matrix is typically inaccessible without custom implementations. Thus, while both methods reuse a precomputed impedance representation, MAPES delivers broader modeling flexibility, easier tool integration, and superior scalability for large pixel design spaces. Finally, MoM requires fine meshing (e.g., RWG basis on triangular elements [28]) to resolve pixel details, inflating the number of unknowns and yielding dense matrices that scale poorly, especially for multilayer arrays, whereas MAPES places ports only at electrically essential locations, producing a far more compact representation and smaller matrices.
@@ IV-C Assumptions and Settings for One-Time Pre-Simulation
IV-D Potential Applications of MAPES

### v2

@@ I Introduction
Secondly, for specific specifications (targeted scattering parameters), the developed efficient EM simulators can be utilized for inverse synthesis of EM structures using heuristic algorithms [5, 15], such as genetic algorithms (GA), simulated annealing, or generative AI tools such as auto-encoders, tandem neural networks [16], or diffusion models [12].
@@
Another technique conceptually related to MAPES, the Method of Moments (MoM)-based "efficient calculator" in [28], precomputes the impedance matrix of a "mother" structure and accelerates exploration by directly manipulating this matrix (removing rows/columns) to evaluate substructures, thereby avoiding repeated matrix refills/factorizations. However, that framework was devised for planar plates and thin-wire geometries excited by voltage/current sources, making it difficult to incorporate vertical ports, vias, and lumped or active components-elements that are intrinsic to MW/RFIC pixels. In contrast, MAPES natively captures horizontal and vertical couplings and supports arbitrary loads, including realistic nonzero via impedances and comprehensive I/O modeling. Furthermore, the MAPES prior-data impedance matrix can be obtained directly from commercial tools (e.g., CST, HFSS), while the mesh-level MoM impedance matrix is typically inaccessible without custom implementations. Thus, while both methods reuse a precomputed impedance representation, MAPES delivers broader modeling flexibility, easier tool integration, and superior scalability for large pixel design spaces. Finally, MoM requires fine meshing (e.g., RWG basis on triangular elements [29]) to resolve pixel details, inflating the number of unknowns and yielding dense matrices that scale poorly, especially for multilayer arrays, whereas MAPES places ports only at electrically essential locations, producing a far more compact representation and smaller matrices. It is worth noting that the recently proposed precomputed numerical Green-function framework [30] likewise absorbs the static environment into a single operator obtained from full-wave solvers, but it operates on voxel- or mesh-level field unknowns, so its system matrices remain substantially larger than those of MAPES and scale with mesh density, similar to the MoM-based efficient calculator.
@@ IV-C Assumptions and Settings for One-Time Pre-Simulation
IV-D Comparison with MoM-Based Efficient Calculator
Recently, the precomputed numerical Green-function (PNGF) framework in [30] extended the MoM-style “efficient calculator” philosophy to general full-wave environments. In PNGF, the static environment is absorbed into a numerical Green matrix $\mathbf{G}$, obtained once using either custom FDTD/FDFD solvers or commercial tools such as HFSS/CST, and candidate layouts are then evaluated by solving reduced linear systems for equivalent polarization currents confined to the optimization region. Compared to the original MoM efficient calculator, PNGF therefore inherits two important advantages that are also characteristic of MAPES: (i) the underlying operator ($\mathbf{G}$) can be generated directly by commercial full-wave solvers without exposing internal MoM matrices, and (ii) fully three-dimensional structures with vertical features, stacked substrates and waveguiding elements can be treated in a unified manner. From the viewpoint of pixel-based RFIC/PCB layout exploration, however, PNGF operates on a system matrix whose dimension is the number of field unknowns in the optimization region, $N_{\mathrm{opt}}$, i.e., the number of Yee-cell edges or FEM/BEM basis functions. In practice $N_{\mathrm{opt}}$ typically scales as $N_{\mathrm{opt}}\approx N_{\mathrm{pixel}}\times N_{\mathrm{cell/pixel}}$, and accurate discretizations often require tens of mesh cells per pixel (e.g., 20–30), so the linear-algebra cost and memory footprint remain substantially larger than those of MAPES, whose core matrices have size $(4\text{--}6)\,\times N_{\mathrm{pixel}}$ and are independent of mesh density. Furthermore, the Woodbury low-rank update strategy [30, 32] used in PNGF to accelerate direct-binary-search optimization when only one or a few tiles change per iteration is equally applicable to MAPES: when a design step modifies only a small subset of pixel/via loads in $\mathbf{Z}_{L}$, the effective I/O matrix in (7) can, in principle, be updated via low-rank corrections rather than recomputing a full inverse, offering an additional route to further speed up large-scale MAPES-based inverse design.
IV-E Potential Applications of MAPES